OpenStudy (anonymous):

Please confirm my answer! Solve for unknown part of triangle: a=22, b=31, c=40. C=96 degrees 30'?

OpenStudy (anonymous):

I see you guys typing but nothing is showing up :/

hero (hero):

Which part was the "unknown" part?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Which I got 96 degrees 30' but I need a confirmation

OpenStudy (eyust707):

There are two unknown angles \[{\sin96 \over 40} = {\sin A \over 22} \] \[22 \sin96 = 40 \sin A\] \[{22 \sin 96 \over 40} = \sin A\] \[\sin^{-1} ({11 \sin96 \over20}) = A\] \[A \approx 33.16 ^{\text o} \]

hero (hero):

So C = 96 degrees, but what is 30'?

OpenStudy (anonymous):

a. 96°30' b. 95°0' c. 85°0' d. 83°30' These are the options

hero (hero):

eyust707, she only needed to find angle C

OpenStudy (eyust707):

I am confused

hero (hero):

Oh, I see what the 30' is. 96 degrees, 30 minutes

OpenStudy (eyust707):

ahh

hero (hero):

 I'm thinking she was given something similar to this